OpenStudy (anonymous):

Which is the most reasonable estimate for 25% of 1,182? a. 250 b. 300 c. 350 d. 400

OpenStudy (anonymous):

b

OpenStudy (anonymous):

295.5

OpenStudy (anonymous):

25% of 1,182\[= \frac{25}{100} \times 1182 = \frac{1}{4} \times 1182 =295.5\]

OpenStudy (anonymous):

Thanks guys!!!

OpenStudy (tkhunny):

Really quickly. 1182 isn't too far from 1200. 1/4 of 1200 is 300. If we're looking for "most reasonable", I think we have it at 300.

OpenStudy (anonymous):

Thanks